# Hybrid second-order gradient histogram based global low-rank sparse regression for robust face recognition

### v1

@@ Abstract
Low-rank sparse regression models have been widely applied in the field of face recognition. To further address the challenges caused by complex occlusions and illumination variations, this paper proposes a Hybrid Second-Order Gradient Histogram based Global Low-Rank Sparse Regression (H2H-GLRSR) model. Specifically, a novel feature descriptor called the Hybrid Second-Order Gradient Histogram (H2H) is first designed to more effectively characterize the local structural features of facial images. Then, this descriptor is integrated with the Sparse Regularized Nuclear Norm based Matrix Regression (SR_\_NMR). Moreover, a global low-rank constraint is imposed on the residual matrix, enabling the model to better capture the global correlations inherent in structured noise. Experimental results demonstrate that the proposed method significantly outperforms existing regression-based classification approaches under challenging scenarios involving occlusions, illumination changes, and unconstrained environments.
@@ 1 Introduction
Although methods such as HSOG have been effective, they often rely on complex feature construction techniques (e.g., DAISY-style ring pooling) to enhance discriminability, which inevitably increases computational cost. Inspired by this, we first propose a histogram of oriented hessian (HOH) directly based on the Hessian matrix, which effectively captures local geometric properties related to curvature by statistically aggregating the principal direction and magnitude of second-order gradients over a uniform grid. Furthermore, we parallelly compute HOG and HOH on the same cell grid and combine them at the block level with normalization, thus constructing a more efficient hybrid second-order gradient histogram (H2H) feature descriptor. This descriptor maintains the complementary advantages of first-order and second-order gradients while significantly simplifying the feature construction process and improving computational efficiency. Additionally, to further enhance the model’s robustness to structured noise and large-area occlusion, we improve the residual modeling process based on the SR_NMR framework and propose a hybrid second-order gradient histograms based global low-rank sparse regression (H2H-GLRSR) model. In contrast to existing methods that impose sample-wise low-rank constraints on the residuals (e.g., sample-wise nuclear norm minimization), our model imposes a global nuclear norm constraint, applying a joint low-rank regularizer to the entire residual matrix, thereby effectively capturing the global correlations among residuals and improving the face recognition performance in complex environments.
@@ 3 The proposed method
In this section, we first introduce the proposed HOH, which captures second-order geometric characteristics such as curvature and ridge structures, and then develop H2H that integrates HOH with first-order gradient features to form a unified representation of local structural information. Subsequently, we incorporate H2H features into a Global Low-Rank Sparse Regression (GLRSR) framework, leading to the H2H-GLRSR model. Furthermore, we present an efficient ADMM-based optimization algorithm for solving the model, followed by a theoretical complexity analysis and a ridge regression–based classification strategy.
@@ 3.3 Hybrid second-order gradient histograms based global low-rank sparse regression (H2H-GLRSR)
To overcome the limitations of SR_NMR in handling structured noise and correlated occlusions across samples, we extend the SR_NMR framework by introducing a global low-rank constraint on the residual matrix. This extension leads to the proposed H2H-GLRSR model, which simultaneously exploits the discriminative power of H2H features and the global correlation structure of residuals.
@@ 3.3.1 Fast optimization algorithm for H2H-GLRSR
$\boldsymbol{Z}^{k+1}=\left(\boldsymbol{X}_{tr}^{\top}\boldsymbol{X}_{tr}+\boldsymbol{I}_{n}\right)^{-1}\left[\boldsymbol{X}_{tr}^{\top}\left(\boldsymbol{Y}-\boldsymbol{E}^{k}+\frac{1}{\mu^{k}}\boldsymbol{T}_{1}^{k}\right)+\boldsymbol{J}^{k}-\frac{1}{\mu^{k}}\boldsymbol{T}_{2}^{k}\right].$ (28)
@@
1. Compute $\boldsymbol{C}:=\boldsymbol{X}_{tr}^{\top}\boldsymbol{X}_{tr}+\boldsymbol{I}_{n}$ and perform LU decomposition on $\boldsymbol{C}$;
@@ 3.4 Computational Complexity Analysis
Similar to [2019A], we neglect the computational cost of simple matrix addition, subtraction, and scalar matrix operations, since they are negligible compared to matrix decomposition. The main computational burden of the proposed algorithm comes from two parts: (1) the update of the regression coefficient matrix $\boldsymbol{Z}$; (2) the low-rank reconstruction of the residual matrix $\boldsymbol{E}$.
In each iteration, instead of directly computing the inverse of $(\boldsymbol{X}_{tr}^{T}\boldsymbol{X}_{tr}+\boldsymbol{I})$ as in[2019A], we perform LU decomposition once at the beginning of the iteration process and reuse it in all subsequent iterations. This reduces the computational cost of each $\boldsymbol{Z}$-update from $O(n^{3}+n^{2}m)$ (for direct matrix inversion) to $O(n^{2}m)$.
In the residual update stage, all test samples are unified into a single residual matrix $\boldsymbol{E}$, and its low-rank structure is recovered via a single singular value decomposition (SVD), the complexity of this SVD step is $O\big(\min\{d^{2}m,dm^{2}\}\big)$. Since only one SVD is required for all test samples in each iteration, our approach avoids the repetitive computations needed for performing SVD on each test image separately as described in [2019A].
Therefore, the overall computational complexity of the proposed method is approximately $O\Big(n^{2}m+t\times\min\{d^{2}m,dm^{2}\}\Big)$,where $t$ is the number of iterations.
@@ 4.2 Face images with illumination changes
Table 2 summarizes the recognition results of different methods under varying illumination conditions on the Extended Yale B database. The proposed H2H-GLRSR method achieves recognition rates of 99.8% and 99.3% on Subset 4 and Subset 5, respectively, outperforming all compared methods. In contrast, traditional sparse or collaborative representation-based models (SRC, CRC) experience severe performance degradation under strong illumination variations (Subset 5), with recognition rates dropping to 28.8% – 43.2%, reflecting their limited robustness. Although the recently proposed ID-NMR method significantly improves recognition performance to 97.4% and 96.9%, demonstrating strong competitiveness, H2H-GLRSR still maintains a consistent advantage across both subsets. This improvement can be attributed to the incorporation of HOH features, which are more sensitive to image curvature and local geometric structure. By statistically analyzing the principal directions and magnitudes of second-order differentials, HOH enhances the discriminability and stability of features under drastic illumination variations. These results confirm the unique value of integrating second-order gradient information for effectively handling illumination changes in face recognition.
@@ 4.3 Face images in unconstrained environment
Figure 4: Some sample images of one person from the LFW database
@@ 4.5 Computation time
The above methods are all implemented in Matlab R2025a. The experiments were conducted on a 64-bit, 3.70 GHz Intel Core Ultra 9285K (multi-core) CPU computer with 192 GB of memory. From the running time comparison results, it can be seen that the proposed H2H-GLRSR method has significant advantages in computational efficiency. Specifically, on the AR dataset, H2H-GLRSR takes only 23.62 seconds in the sunglasses case, which is about 11 times and 134 times faster than SR_NMR and ID-NMR, respectively; in the scarf case, it takes 69.00 seconds, which is about 3.7 times and 46 times faster. On the Extended Yale dataset, H2H-GLRSR takes 110.53 seconds and 209.56 seconds in the subset4 and subset5 cases, respectively, which is 1.4 to 4.3 times faster than the comparison methods. In the unconstrained LFW environment, H2H-GLRSR takes only 8.27 seconds, showing the most prominent efficiency improvement, which is about 22 times and 44 times faster than SR_NMR and ID-NMR, respectively.

### v2

@@ Abstract
Low-rank sparse regression models have been widely adopted in face recognition due to their robustness against occlusion and illumination variations. However, existing methods often suffer from insufficient feature representation and limited modeling of structured corruption across samples. To address these issues, this paper proposes a Hybrid second-order gradient Histogram based Global Low-Rank Sparse Regression (H2H-GLRSR) model. First, we propose the Histogram of Oriented Hessian (HOH) to capture second-order geometric characteristics such as curvature and ridge patterns. By fusing HOH and first-order gradient histograms, we construct a unified local descriptor, termed the Hybrid second-order gradient Histogram (H2H), which enhances structural discriminability under challenging conditions. Subsequently, the H2H features are incorporated into an extended version of the Sparse Regularized Nuclear Norm based Matrix Regression (SR_NMR) model, where a global low-rank constraint is imposed on the residual matrix to exploit cross-sample correlations in structured noise. The resulting H2H-GLRSR model achieves superior discrimination and robustness. Experimental results on benchmark datasets demonstrate that the proposed method significantly outperforms state-of-the-art regression-based classifiers in both recognition accuracy and computational efficiency.
@@ 1 Introduction
Although methods such as HSOG have been effective, they often rely on complex feature construction techniques (e.g., DAISY-style ring pooling) to enhance discriminability, which inevitably increases computational cost. Inspired by this, we first propose a histogram of oriented hessian (HOH) directly based on the Hessian matrix, which effectively captures local geometric properties related to curvature by statistically aggregating the principal direction and magnitude of second-order gradients over a uniform grid. Furthermore, HOG and HOH features are computed in parallel on the same cell grid and combined at the block level with normalization, resulting in an efficient hybrid second-order gradient histogram (H2H) feature descriptor. This descriptor retains the complementary advantages of first- and second-order gradients while simplifying the feature construction process and improving computational efficiency. To further enhance the model’s robustness to structured noise and large-area occlusion, the H2H features are incorporated into the SR_NMR framework, and the residual modeling process is augmented with a global low-rank constraint. This integration gives rise to the hybrid second-order gradient histogram based global low-rank sparse regression (H2H-GLRSR) model. In contrast to existing methods that impose sample-wise low-rank constraints on the residuals (e.g., sample-wise nuclear norm minimization), our model imposes a global nuclear norm constraint, applying a joint low-rank regularizer to the entire residual matrix, thereby effectively capturing the global correlations among residuals and improving the face recognition performance in complex environments.
@@ 3 The proposed method
In this section, we first introduce the proposed HOH, which captures second-order geometric characteristics such as curvature and ridge-like structures. We then develop the H2H by integrating HOH with first-order gradient features, forming a unified representation of local structural information. Subsequently, the H2H features are incorporated into the SR_NMR model. By imposing a global low-rank constraint on the residual matrix, the resulting model, referred to as H2H-GLRSR, effectively exploits cross-sample correlations in structured noise while maintaining sparsity. Furthermore, we use an efficient ADMM-based optimization algorithm for solving the model, followed by a theoretical complexity analysis and a ridge regression–based classification strategy.
@@ 3.3 Hybrid second-order gradient histograms based global low-rank sparse regression (H2H-GLRSR)
To overcome the limitations of SR_NMR in handling structured noise and correlated occlusions across samples, a global low-rank constraint is imposed on the residual matrix. Incorporating this constraint with the H2H features gives rise to the H2H-GLRSR model, which simultaneously leverages the discriminative power of H2H features and captures the global correlation structure of the residuals.
@@ 3.3.1 Fast optimization algorithm for H2H-GLRSR
$\boldsymbol{Z}^{k+1}=\boldsymbol{C}\left[\boldsymbol{X}_{tr}^{\top}\left(\boldsymbol{Y}-\boldsymbol{E}^{k}+\frac{1}{\mu^{k}}\boldsymbol{T}_{1}^{k}\right)+\boldsymbol{J}^{k}-\frac{1}{\mu^{k}}\boldsymbol{T}_{2}^{k}\right],$ (28)
where $\boldsymbol{C}=(\boldsymbol{X}_{tr}^{\top}\boldsymbol{X}_{tr}+\boldsymbol{I}_{n})^{-1}$.
@@
1. Compute $\boldsymbol{C}=(\boldsymbol{X}_{tr}^{\top}\boldsymbol{X}_{tr}+\boldsymbol{I}_{n})^{-1}$;
@@ 3.4 Computational Complexity Analysis
Similar to [2019A], we neglect the computational cost of simple matrix addition, subtraction, and scalar matrix operations, since they are negligible compared to matrix decomposition. The main computational burden of the proposed algorithm comes from two parts: (1) calculating the inverse of matrix in (28); (2) the low-rank reconstruction of the residual matrix $\boldsymbol{E}$.
Before the iteration begins, the complexity of pre-computing $(\boldsymbol{X}_{tr}^{\top}\boldsymbol{X}_{tr}+\boldsymbol{I}_{n})^{-1}$ in (28) is $O(n^{3})$. In the residual update stage, all test samples are unified into a single residual matrix $\boldsymbol{E}$, and its low-rank structure is recovered via a single singular value decomposition (SVD), the complexity of this SVD step is $O\big(\min\{d^{2}m,dm^{2}\}\big)$. Since only one SVD is required for all test samples in each iteration, our approach avoids the repetitive computations needed for performing SVD on each test image separately as described in [2019A].
Therefore, the overall computational complexity of the proposed method is approximately $O\Big(n^{3}+t\times\min\{d^{2}m,dm^{2}\}\Big)$,where $t$ is the number of iterations.
@@ 4.2 Face images with illumination changes
Table 2 reports the recognition accuracies of various methods under different illumination conditions on the Extended Yale B database. Traditional sparse or collaborative representation models (SRC, CRC, RSC) suffer dramatic performance degradation under extreme illumination (Subset 5), where their accuracies decrease to approximately 28.8% – 36.7%, indicating limited robustness to lighting variations. More advanced low-rank or regularized models (RLRC, RNR, NMR, SR_NMR) exhibit improved resistance to illumination but still experience noticeable drops when moving from Subset 4 to Subset 5. Although the recently proposed ID-NMR achieves competitive results of 98.4% and 99.4% on the two subsets, respectively, H2H-GLRSR maintains superior performance on Subset 4 (99.8%) and remains nearly comparable on Subset 5 (99.3%), demonstrating strong overall robustness across illumination conditions. The performance gain primarily stems from the introduction of HOH features, which capture curvature-related patterns and second-order geometric structures that are more resilient to drastic lighting changes. By exploiting the principal directions and magnitudes of second-order differentials, HOH enhances feature discriminability and stability. These results validate the effectiveness and unique advantages of incorporating second-order gradient information for illumination-robust face recognition.
@@ 4.3 Face images in unconstrained environment
Figure 4: Some sample images of one person from the LFW database
@@ 4.5 Computation time
The above methods are all implemented in Matlab R2024b. The computational experiments were carried out on a 64-bit workstation equipped with an Intel Core i9-13900K processor operating at 3.00 GHz, 96 GB of system memory, and an Intel UHD Graphics 770 integrated GPU. The machine was configured with approximately 11.85 TB of available storage.
The experimental results on running time, as summarized in Table 4, demonstrate the superior computational efficiency of the proposed H2H-GLRSR method. H2H-GLRSR achieves significantly faster processing speeds across all tested datasets (AR, Extended Yale, and LFW) compared to the competing methods SR_NMR and ID-NMR. Specifically, it runs approximately 80 times faster than ID-NMR and nearly 10 times faster than SR_NMR on the AR dataset, while maintaining a substantial speed advantage (1.7x to 4.5x faster than competitors) on the Extended Yale subsets. Most notably, on the LFW dataset, H2H-GLRSR completes the task in only 9.01 seconds, which is over 19 times and 30 times faster than SR_NMR and ID-NMR, respectively.